Consider the following cost minimization problem:

Minimization CT = wL + rK

Hold it: q= L1/2 K1/2

a) Obtain the long-run cost function

b) If K = 9 and w = 1, obtain the long-run cost function


1
Expert's answer
2021-10-14T10:49:42-0400
"\frac{\delta q}{\delta L}=\frac{1}{2}\times(\frac{K}{L})^{0.5}"

"\frac{\delta q}{\delta K}=\frac{1}{2}\times (\frac{L}{K})^{0.5}"

"r(\frac{K}{L})^{0.5}=w(\frac{L}{K})^{0.5}"

"K=\frac{wL}{r}"

If K = 9 and w = 1


"L=9r"

"CT=9r+9r=18r"